Suppose that the demand and supply curves for a good are given by QD = 50 – P and QS = 4P – 30. At what price is there an excess demand of 40 units?
Answer #1

for excess demand there is Shortage → Shortage a demand - Supply 402 D-S = Sorp - (Up-30) = 50-p-up +30 40 = 80-5p sp= 40 - 4